# Quantum Down Sampling Filter for Variational Auto-encoder

### v2

@@ 3.3.2 Transposed Convolutions
Once the latent variable is transformed into the higher-resolution space, it is passed through a series of transposed convolution (also known as de-convolution) layers.These layers up-sample the tensor, progressively restoring the original spatial dimensions of the input data. Each transposed convolution layer learns to refine the spatial resolution of the image while preserving important features. Transposed convolution layers have been shown to be highly effective in generating high-resolution outputs from low-resolution latent representations. They are capable of learning hierarchical features of the data, capturing both local details and global structure during the reconstruction process. By applying non-linear activation functions, such as ReLU, between the transposed convolution layers, the model introduces the necessary complexity to capture intricate patterns in the data.
@@ 3.3.3 Reconstruction
The final stage of the decoding process involves producing the reconstructed image. This is typically done by passing the output of the last transposed convolution layer through a sigmoid activation function to ensure that the pixel values lie within the valid range (e.g., between 0 and 1 for normalized image data). The output of this step is the reconstructed image, which should ideally match the original input data.
3.4 Quantum Computing in Variational Auto-encoder
Quantum computing has shown promise in enhancing generative models due to its ability to represent information in quantum states. Unlike classical models, which rely on deterministic mappings, quantum encoders can utilize superposition and entanglement to encode data into a potentially much larger computational space, enabling the model to learn richer and more compact latent representations [20]. Quantum circuits can efficiently explore large state spaces, leveraging the Hilbert space to represent data points in ways that are not possible with classical methods [15]. Mathematically, quantum encoders can represent the latent variables $\mathbf{z}$ as quantum states $|\psi\rangle\in\mathcal{H}$, where $\mathcal{H}$ is the Hilbert space. A quantum encoder might take the form of a unitary operator $U_{\theta}$ applied to an initial quantum state $|\mathbf{x}\rangle$, mapping the input data $\mathbf{x}$ to a quantum state that encodes the latent representation that map the medium-dimensional data into a more compact, or lower-dimensional space.
$|\psi\rangle=U_{\theta}|\mathbf{x}\rangle,$ (4)
where $U_{\theta}$ is a quantum operation that performs the encoding. This encoding could be leveraged to provide a quantum-enhanced latent space that could more efficiently represent complex structures within the data. Quantum states encoded in this way can capture correlations in data that are hard to model classically, such as in the case of medium-resolution images.
The advantage of quantum properties has potential to improve tasks like image reconstruction and anomaly detection, overcoming challenges like blurred images often encountered with classical VAEs.
@@ 4 Proposed Methodology
In this work, we introduce a novel down-sampling filter for quantum variational auto-encoders (Q-VAE) that relies exclusively on quantum encoding for feature extraction. The paper investigates three distinct architectures for Variational Auto-encoders (VAEs), each employing different encoding mechanisms to evaluate their feature extraction capabilities: Quantum VAE (Q-VAE), Classical VAE (C-VAE), and Direct Passing VAE (CDP-VAE). The framework of these models has been presented in Figure 1. The goal of this comparison is to assess the effectiveness of quantum computing, classical neural networks, and simplified encoding strategies in capturing and processing data for subsequent reconstruction. While each model shares a common decoder architecture, the encoders differ in their approach, enabling the study of how each influences the encoding process and the quality of the reconstructed output.
@@ 4.1 Classical VAE (C-VAE) Encoder
The Classical VAE (C-VAE) uses a traditional fully connected neural network as the encoder, which processes the input data through a series of hidden layers to output the mean ($\mu$) and log variance ($\log(\sigma^{2})$) of the latent space distribution. This distribution is then sampled to obtain the latent variable $z$, which represents the encoded data in a lower-resolution space. While the C-VAE relies entirely on classical neural networks for encoding, it remains highly effective in modeling data distributions and is widely used in practical applications. The decoder, typically a CNN, reconstructs the input from this low-resolution latent space, generating an output image or data structure that approximates the original input, also shown in Figure 1.
@@ 4.2 Classical Direct Passing (CDP-VAE) Encoder
The Direct Passing VAE (CDP-VAE) adopts a more simplified approach to encoding by using windowing pooling. In this method, the input data (e.g., an image) is divided into 2x2 pixel windows, and only the first pixel of each window is retained, discarding the other three. This process reduces the resolution of the input data while preserving the most significant features. The reduced input is then passed through a fully connected layer for further encoding into the latent space. The CDP-VAE encoder is particularly useful for computationally constrained environments, where reducing the complexity of the encoding process can lead to faster computations while still capturing essential information for reconstruction. The Classical Direct Passing Architecture (CDP) for the encoder as shown in Figure 1 introduces a method to reduce resolution and simplify data processing. Instead of fully processing all pixels through multiple transformations, this architecture selectively passes pixel values directly to subsequent layers. Specifically, it utilizes a windowing approach where a group of four neighboring pixels is considered, but only the first pixel in the window is retained, while the remaining three pixels are discarded. This selective passing reduces the input resolution, focusing on preserving key information. The retained pixel is then forwarded through an fully connected layer for further processing.
Figure 2: USPS image Direct Passing filter conversion to 16x16
Windowing and Pooling for Resolution Reduction: The input image $x$ undergoes a selective passing mechanism, where only the first pixel from every 2x2 window is retained. This operation can be mathematically represented as:
$[x_{\text{reduced}}=\text{Resample}(x),$ (5)
where $x_{\text{reduced}}$ represents the resolution-reduced image also shown in figure 2, and $\text{Resample}(x)$ is a function that selects every other pixel from the original image to achieve the desired down-sampling.
For instance, if the input image $x$ is of dimension $32\times 32$, this resolution reduction process reduces it to $16\times 16$.
@@ 4.3 Quantum VAE (Q-VAE) Encoder
The Quantum VAE (Q-VAE) integrates quantum computing into the encoding process, replacing the traditional classical neural network encoder with a quantum circuit as shown in Figure 1. The quantum encoder maps the input data into a quantum state representation, leveraging quantum superposition to capture complex relationships in the data. Quantum encoding enhances performance by converting the input into quantum states, enabling the Q-VAE to apply rotations to the data. The quantum-enhanced feature extraction is paired with a classical CNN decoder that reconstructs the input data, combining the strengths of quantum processing for feature extraction with classical techniques for data reconstruction.
The Q-VAE encoder compresses input data by encoding it into a lower-resolution latent space representation. It consists of a quantum circuit as shown in Figure 3, integrated with a classical fully connected layer.
The encoder transforms the input data $X$ into two key outputs: the mean $\mu$ and the log variance $\log(\sigma^{2})$ of the latent space distribution:
Figure 3: Quantum circuit for Quantum Variational Auto-encoder
@@
The Pauli-$Z$ operator is applied to measure the quantum state and extract the latent variables $z$. The Pauli-$Z$ measurement flips the qubit’s state depending on the computational basis, effectively capturing the binary features of the image as shown in Figure 4.
Figure 4: USPS Image resolution reduction from Quantum Circuit
Mathematically, the Pauli-$Z$ operator acts on the quantum state $\left|\psi\right\rangle$ as follows:
$Z\left|0\right\rangle=\left|0\right\rangle,\quad Z\left|1\right\rangle=-\left|1\right\rangle$ (10)
where $Z$ flips the phase of the qubit in the $\left|1\right\rangle$ state, while leaving the $\left|0\right\rangle$ state unchanged. This operation effectively translates quantum information into meaningful latent variables for the model. The resulting latent variables $z$ are given by:
@@
After the quantum circuit processes the input image and extracts quantum features, the output of the quantum circuit, as shown in figure 4, is passed to fully connected layer. This transition enables the integration of quantum-encoded features with classical deep learning architectures. The quantum circuit produces an output, typically in the form of a reduced-resolution feature map or a vector. Let the output from the quantum circuit be denoted as $q_{\text{out}}$, where:
$q_{\text{out}}=\text{QuantumCircuit}(x)$ (12)
Here, $x$ is the input image, and $q_{\text{out}}$ represents the quantum-enhanced feature representation. The resolutionity of $q_{\text{out}}$ corresponds to the number of measurements or qubits used in the quantum circuit.
@@ 4.4 Shared Decoder Architecture
Despite the differences in their encoding mechanisms, all three models, C-VAE, CDP-VAE and Q-VAE share a common decoder architecture. This decoder is responsible for reconstructing the input data from the latent space representation. The decoder begins by transforming the latent variable \( z \) into a medium-resolution space through fully connected layers. Then, a series of transposed convolution (de-convolution) layers are used to progressively up-sample the data, restoring the original spatial dimensions. Finally, the output is passed through a sigmoid activation function to ensure that the pixel values are within the valid range (e.g,, between 0 and 1 for image data). This shared decoder allows for a fair comparison of the encoding techniques, with the focus on how the different encoders influence the quality of the reconstructed image.
@@ 5 Numerical Simulation
To assess the performance of our proposed model, we first need to preprocess the input image data. The preprocessing steps are as follows:
@@ 5.1 MNIST Dataset Pre-Processing
We utilized MNIST to determine whether the performance improvements observed in USPS could be replicated with MNIST. To enhance the model’s ability to generalize across different image sizes, we perform the following preprocessing steps also shown in Figure 5:
@@
    Resizing to $16\times 16$: After cropping, the image is resized to $16\times 16$ pixels. This step reduces the image resolution, making it easier for the model to process while still retaining key features of the digit. This process will help us to get same resolution dataset like USPS to see if MNIST also show improvement or not.
@@
    Resizing to $32\times 32$: Finally, the image is resized to $32\times 32$ pixels. This larger size is used for feeding the images into the neural network model, which is optimized to process medium resolution images. This resizing allows the model to capture finer details of the digits, enhancing its ability to reconstruct and classify the images effectively[16].
@@
This resizing step ensures compatibility with the input size of the model while retaining the features necessary for effective representation learning.
5.2 USPS Dataset Pre-Processing
Figure 6: USPS Image transformed from 16x16 to 32x32.
USPS dataset has an image size 16x16. The interpolation method used during resizing adjusts the pixel values to fit the larger image grid, enabling the model to be tested with standard image input sizes, such as 32x32. USPS image conversion is shown in Figure 6.
@@ 5.3 Experimental Setup
In this section, we describe the hyper-parameters and settings used for training the models. We are implementing our algorithm on the Pennylane quantum simulator. Hyper-parameters to run the algorithm were carefully selected to ensure stable training while optimizing performance for all the models.
@@ 5.3.1 Hyper-parameters
In our training process, several key hyper-parameters were used to ensure efficient learning and model convergence. The learning rate was set to $0.001$ for all the models. This learning rate controls how much the model’s weights are updated in response to the computed gradients, providing a stable training process while allowing sufficient progress over the course of the training. The models were trained for 200 epochs, with each epoch representing one full pass of the data set through the model. This duration was chosen to ensure that the models had enough time to learn the underlying features of the data, while regular monitoring of the validation loss helped prevent over-fitting. A batch size of $400$ was used, which means that $400$ images were processed together before updating the model parameters. This batch size was selected to balance computational efficiency and the stability of gradient estimation, helping smooth gradient updates and ensuring more stable training. For optimization, we used the Adam optimizer, which is known for its adaptive learning rate that adjusts based on the first and second moments of the gradients. The Adam optimizer is particularly effective for models with many parameters and large datasets. The loss function used was the Binary Cross-Entropy Loss (BCELoss), with the argument reduction = ’sum’, which ensures that the individual losses across all pixels are summed rather than averaged. This choice emphasizes the total error in the reconstruction of the images. To visualize the progress of the model over time, the variables were used to store a grid of reconstructed images at various epochs. These images allowed us to visually assess how well the models were reconstructing the images as training progressed. These configurations were chosen to ensure effective model training, allowing the models to learn from the data efficiently and perform well in both reconstruction and evaluation tasks.
6 FID and MSE Performance Evaluation and Discussion
To evaluate the performance of the proposed models, we used a combination of several key metrics: the Fréchet Inception Distance (FID) score[18], training and testing loss, and Mean Squared Error (MSE)[19].
@@ 6 FID and MSE Performance Evaluation and Discussion
Each of these metrics provides distinct insights into the model’s ability to generate better-quality images and generalize to new data: Fréchet Inception Distance (FID) Score measures the similarity between the distribution of real and generated images. A lower FID score indicates that the generated images are closer in distribution to the real images, reflecting higher image fidelity and quality. The FID score is particularly useful for assessing the realism of generated images in generative models. Mean Squared Error (MSE) measures the average squared difference between the generated and real images. A lower MSE indicates that the generated images are closer to the real images in terms of pixel-wise accuracy. This metric is commonly used to evaluate the reconstruction quality in generative models, with lower values indicating better performance. By combining these metrics, we gain a comprehensive understanding of how well our models perform in terms of both image quality and generalizability.
The FID score was computed for 1) images reconstructed from the input images and 2) images generated from random noise, giving us insight into how well the models perform in generating realistic images from these two different sources of input.
The results are shown in table 1 below:
@@
The performance of all the models on the MNIST dataset was assessed by calculating the FID score using latent space representations as shown in Figure 7:
Figure 7: (a) MNIST FID Score for Quantum Variational Auto-encoder and (b) MNIST Direct Passing FID Score for Variational Auto-encoder.
@@
For MNIST, the C-VAE achieves an FID score of 40.7 , indicating a reasonable ability to reconstruct images. CDP-VAE, demonstrates a slight improvement over the C-VAE with an FID score of 39.7 and then to Q-VAE, further improving upon CDP-VAE. The Q-VAE leverages quantum gates such as RY rotations to encode the data into a quantum state and Pauli-Z measurements bring quantum state to classical data, achieving a superior FID score of 37.3. This improved score suggests that the quantum model produces more faithful image reconstructions, capturing complex relationships in the data more effectively than the classical model. Although CDP-VAE approach is less complex, it still lacks the advanced feature extraction capabilities of the Q-VAE, which ultimately leads to lower reconstruction fidelity.
@@
For the USPS dataset, the Q-VAE achieved an impressive FID score of 57.6, indicating better-quality image reconstructions and strong performance in terms of both fidelity and distribution similarity to the original dataset. This result highlights the effectiveness of the Q-VAE in handling simpler grayscale datasets, where it demonstrates superior feature extraction and reconstruction capabilities.
Figure 8: (a) USPS FID Score for Variational Quantum Auto encoder and (b) USPS Direct Passing FID Score for Variational Auto encoder.
Despite its simplicity, the CDP-VAE still achieved relatively good performance on the USPS dataset, with an FID score of 66.1 as shown in Figure 8.
On the MNIST dataset, C-VAE, with an MSE of 0.0093, showed slightly higher error compared to the Q-VAE, indicating that while it performs well, its feature extraction capabilities are not as advanced as those of the quantum model. Despite its relatively low MSE, the classical model’s performance is constrained by the limitations of traditional neural network-based encoding and decoding methods, which cannot capture the full complexity of the data as efficiently as quantum circuits. The CDP-VAE, with an MSE of 0.0092, performs little better then C-VAE While Q-VAE achieved the lowest MSE of 0.0088, demonstrating its ability to produce high-fidelity reconstructions compared to the other models. This is consistent with the Q-VAE’s superior performance across other metrics such as training and testing loss. The quantum-based encoding method enables the Q-VAE to extract richer feature representations, leading to more accurate reconstructions. The quantum model’s ability to capture complex dependencies between pixels allows it to outperform classical approaches, even when dealing with highly structured images like those in the MNIST dataset.
6.1 Reconstructed Images:

### v3

@@ 3.4.2 Transposed Convolutions
Once the latent variable is transformed into the higher-dimensional space, it is passed through a series of transposed convolution layers. These layers up-sample the tensor, progressively restoring the original spatial dimensions of the input data. Each transposed convolution layer learns to refine the spatial resolution of the image while preserving important features. Transposed convolution layers have been shown to be highly effective in generating high-resolution output from low-resolution latent representations. They are capable of learning hierarchical features of the data, capturing both local details and global structure during the reconstruction process. By applying non-linear activation functions, such as ReLU, between the transposed convolution layers, the model introduces the necessary complexity to capture intricate patterns in the data.
@@ 3.4.3 Reconstruction
The final stage of the decoding process involves the production of the reconstructed image. This is typically done by passing the output of the last transposed convolution layer through a sigmoid activation function to ensure that the pixel values lie within the valid range (e.g., between 0 and 1 for normalized image data). The output of this step is the reconstructed image, which should ideally match the original input data.
@@ 4 Proposed Methodology
In this work, we introduce a novel down-sampling filter for Q-VAE that relies exclusively on quantum encoding for feature extraction. The paper investigates three distinct architectures for VAEs, each employing different encoding mechanisms to evaluate their feature extraction capabilities: C-VAE, CDP-VAE and Q-VAE. The framework of these models is presented in Figure 1. The goal of this comparison is to assess the effectiveness of quantum computing, classical neural networks, and simplified encoding strategies in the capture and processing of data for subsequent reconstruction. Although each model shares a common decoder architecture, the encoders differ in their approach, enabling the study of how each influences the encoding process and the quality of the reconstructed output.
@@ 4.1 Classical VAE (C-VAE) Encoder
C-VAE uses a traditional fully connected neural network as the encoder, which processes the input data through a series of hidden layers to output the mean ($\mu$) and log variance ($\log(\sigma^{2})$) of the latent space distribution. This distribution is then sampled to obtain the latent variable $z$, which represents the encoded data in a lower-resolution space. While the C-VAE relies entirely on classical neural networks for encoding, it remains highly effective in modeling data distributions and is widely used in practical applications. The decoder,which employs layers for up-sampling, reconstructs the input from this low-dimension latent space, generating an output image or data structure that approximates the original input, also shown in Figure 1 (a).
@@ 4.2 Classical Direct Passing (CDP-VAE) Encoder
CDP-VAE adopts a more simplified approach to encoding by using windowing pooling for resolution reduction as shown in Figure 1 (b). In this method, the input data (e.g., an image) is divided into $2\times 2$ pixel windows, and only the first pixel of each window is retained, discarding the other three. This process reduces the resolution of the input data while preserving significant features. The reduced input is then passed through a fully connected layer for further encoding into the latent space. The CDP-VAE encoder is particularly useful for computationally constrained environments, where reducing the complexity of the encoding process can lead to faster computations while still capturing essential information for reconstruction. For instance, if the input image $x$ is of dimension $32\times 32$, this resolution reduction process reduces it to $16\times 16$.
@@ 4.3 Quantum VAE (Q-VAE) Encoder
Q-VAE integrates quantum computing into the encoding process, adding in the CDP-VAE as shown in Figure 1 (c). The quantum encoder processes input data into a quantum state representation, utilizing quantum superposition to represent intricate relationships within the data. Quantum encoding improves performance by transforming input data into quantum states, allowing the Q-VAE to apply rotations that facilitate efficient processing and capture richer, more intricate representations. The quantum-enhanced feature extraction is combined with a classical decoder that uses transposed convolution to reconstruct the input data, combining the strengths of quantum processing for feature extraction with classical techniques for data reconstruction.
Figure 2: Proposed Quantum circuit for Quantum Variational Auto-encoder
@@
The Pauli-$Z$ operator is applied to measure the quantum state. The Pauli-$Z$ measurement flips the qubit’s state depending on the computational basis, effectively capturing the binary features of the image.
Mathematically, the Pauli-$Z$ operator acts on the quantum state $\left|\psi\right\rangle$ as follows: where $Z$ flips the phase of the qubit in the $\left|1\right\rangle$ state, while leaving the $\left|0\right\rangle$ state unchanged. This operation effectively translates quantum information into meaningful latent variables for the model. The resulting latent variables $z$ are given by:
@@
The measurement on qubit yields the compressed representation of the input. This procedure reduces the dimension of the input image from 1024 to 256 measured values. After the quantum circuit processes the input image and extracts quantum features, the output of the quantum circuit, as shown in Figure 1 (c), is passed to fully connected layer. This transition enables the integration of quantum-encoded features with classical deep learning architectures.
@@ 4.4 Shared Decoder Architecture
Despite the differences in their encoding mechanisms, all three models, Q-VAE, C-VAE, and CDP-VAE, share a common decoder architecture. This decoder is responsible for reconstructing the input data from the latent space representation. The decoder begins by transforming the latent variable (z) into a medium-resolution space through fully connected layers. Then a series of transposed convolution (deconvolution) layers are used to progressively up-sample the data, restoring the original spatial dimensions. Finally, the output is passed through a sigmoid activation function to ensure that the pixel values are within the valid range (e.g,, between 0 and 1 for image data). This shared decoder allows for a fair comparison of the encoding techniques, with the focus on how the different encoders influence the quality of the reconstructed image.
@@ 5 Numerical Simulation
To assess the performance of our proposed model, we first need to preprocess the input image data. The preprocessing steps are described in the following sections.
5.1 USPS Dataset Pre-Processing
USPS dataset input image size resolution is 16x16. The USPS images are resized to $32\times 32$ using the same resizing function as was done for MNIST.
Figure 3: USPS Image transformed from 16x16 to 32x32.
The interpolation method used during resizing adjusts the pixel values to fit the larger image grid, enabling the model to be tested with standard image input sizes, such as 32x32. USPS image conversion is shown in Figure 3.
@@ 5.2 MNIST Dataset Pre-Processing
As observed in USPS, We utilized MNIST to determine whether the performance improvements observed in USPS could be replicated with MNIST. To enhance the model’s ability to generalize across different image sizes, we perform the following preprocessing steps:
@@
    Resizing to $16\times 16$: After cropping, the image is resized to $16\times 16$ pixels. This step reduces the image resolution, making it easier for the model to process while still retaining key features of the digit. The resizing step uses an interpolation technique that preserves the overall structure of the digit. This process will help us to get same resolution dataset like USPS to see if MNIST also show improvement or not.
@@
    Resizing to $32\times 32$: Finally, the image is resized to $32\times 32$ pixels. This larger size is used for feeding the images into the neural network model, which is optimized to process medium resolution images.This study investigates how resizing adjusts the image resolution and its impact on the model’s ability to process finer details for reconstruction and classification.
Figure 4 shows the transformation of MNIST images as processed according to the above procedure. This resizing step ensures compatibility with the input size of the model while retaining the features necessary for effective representation learning.
@@ 5.3 Experimental Setup
In this section, we describe the hyper-parameters and settings used for training the models. We are implementing our algorithm on Pennylane quantum simulator. The code is available at https://github.com/RRFar/Quantum-Downsampling-Filter. Hyper-parameters to run the algorithm were carefully selected to ensure stable training while optimizing performance for all models.
@@ 5.3.1 Hyper-parameters
In our training process, several key hyper-parameters were used to ensure efficient learning and model convergence. The learning rate was set to $0.001$ for all the models. This learning rate controls how much the model’s weights are updated in response to the computed gradients, providing a stable training process while allowing sufficient progress over the course of the training. The models were trained for 200 epochs, with each epoch representing one full pass of the data set through the model. This duration was chosen to ensure that the models had enough time to learn the underlying features of the data, while regular monitoring of the validation loss helped prevent over-fitting. A batch size of $400$ was used, which means that $400$ images were processed together before updating the model parameters. This batch size was selected to balance computational efficiency and the stability of gradient estimation, helping smooth gradient updates and ensuring more stable training. For optimization, we used the Adam optimizer, which is known for its adaptive learning rate that adjusts based on the first and second moments of the gradients. The Adam optimizer is particularly effective for models with many parameters and large datasets. The loss function used was the binary cross-entropy loss (BCELoss), which ensures that the individual losses across all pixels are summed rather than averaged. This choice emphasizes the total error in the reconstruction of the images. These images allowed us to visually assess how well the models were reconstructing the images as training progressed. These configurations were chosen to ensure effective model training, allowing the models to learn from the data efficiently and perform well in both reconstruction and evaluation tasks.
6 FID and MSE Score Performance Evaluation and Discussion
@@
To evaluate the performance of the proposed models, we used a combination of several key metrics: FID score [12], training and testing loss, and MSE [13]. Each of these metrics provides distinct insights into the model’s ability to generate better-quality images and generalize to new data: FID measures the similarity between the distribution of real and generated images. A lower FID score indicates that the generated images are closer in distribution to the real images, reflecting higher image fidelity and quality. The FID score is particularly useful for assessing the realism of generated images in generative models. MSE measures the average squared difference between the generated and real images. A lower MSE indicates that the generated images are closer to the real images in terms of pixel-wise accuracy. This metric is commonly used to evaluate the reconstruction quality in generative models, with lower values indicating better performance. MNIST and USPS MSE results are shown in Figure 5 and Figure 6 respectively.
Figure 5: MNIST MSE score for all models
Figure 6: USPS MSE score for all models
By combining these metrics, we gain a comprehensive understanding of how well our models are performing in terms of both image quality and generalization ability. The results are shown in Table 1 below:
@@
The FID score was computed for 1) images reconstructed from the input images and 2) images generated from random noise, giving us insight into how well the models perform in generating realistic images. The variation of FID score as a function of the number of epochs for MNIST is shown in Figure 7.
Figure 7: MNIST FID score for image reconstruction and image generation.
@@
For MNIST, the C-VAE achieves an FID score of 40.7, indicating a reasonable ability to reconstruct images. The Q-VAE, on the other hand, leverages quantum gates such as RY rotations and Pauli-Z measurements to encode the data, achieving a superior FID score of 37.3. This improved score suggests that the quantum model produces more faithful image reconstructions, capturing complex relationships in the data more effectively than the classical model. Finally, the CDP-VAE, demonstrates a slight improvement over the C-VAE with an FID score of 39.7. Similar trends have been noticed for USPS where C-VAE achieved 50.4, CDP-VAE scored 42.9 and Q-VAE FID score for reconstruction as 38.5. Although this approach is less complex, it still lacks the advanced feature extraction capabilities of the Q-VAE, which ultimately leads to lower reconstruction fidelity.
@@
For the USPS dataset image generation, the Q-VAE achieved an FID score of 57.6, while MNIST achieved 78.7 indicating better-quality image reconstructions and strong performance in terms of both fidelity and distribution similarity to the original dataset. While for MNIST C-VAE and CDP-VAE scored 94.4 and 93.3 respectively. Similar trends have been seen for USPS where C-VAE and CDP-VAE scored 73.1 and 66.1 respectively. This result highlights the effectiveness of the Q-VAE in handling simpler grayscale datasets, where it demonstrates feature extraction and reconstruction capabilities.
Figure 8: USPS FID score for image reconstruction and image generation.
The CDP-VAE is a VAE where, instead of using an encoder-decoder architecture with learned transformations, the input data is directly passed through with minimal transformation using windowing/ pooling. Despite its simplicity, the CDP-VAE still achieved relatively good performance on the USPS dataset, with an FID score of 66.1 as shown in Table 1. Figure 8 further illustrates that the Q-VAE consistently achieves a lower FID score for reconstruction compared to both CDP-VAE and C-VAE, indicating its superior ability to generate high-fidelity reconstructions.
6.1 Latent Space Visualization
To better understand the expressiveness of the model, the latent vector mapping is visualized using dimensionality reduction techniques such as t-SNE [37], which shows the clustering structure in the latent space. Furthermore, Gaussian Mixture Model (GMM) clustering is used to assess how well the learnt representations can be grouped, providing information on the model’s capacity to capture meaningful variations and structure in the data. The colours in the visualisation reflect different classes, with well-separated colours indicating efficient feature learning and overlapping colours indicating class similarities or ambiguity. Distinct clusters suggest that the model has learnt meaningful feature representations, but overlapping clusters could indicate insufficient separation in the latent space. By combining t-SNE and GMM clustering, we can assess whether the model effectively organises data points and captures key properties for classification or reconstruction. C-VAE, CDP-VAE and Q-VAE latent space has been visulatise in Figure  910 and  11 respectively.
Figure 9: Latent space for C-VAE
Figure 10: Latent space for CDP-VAE
Figure 11: Latent space for Q-VAE
From the figures, it is evident that CDP-VAE and Q-VAE do not show a significant difference from C-VAE, but demonstrate competitive performance. For further investigation, we will analyze the reconstructed images and newly generated images in the next section.
6.2 Reconstructed Images